# When the Correct Model Fails: The Optimality of Stackelberg Equilibria with Follower Intention Updates

### v1

@@ Abstract
We study a two-player dynamic Stackelberg game between a leader and a follower. Classical formulations of the Stackelberg equilibrium (SE) assume that the follower’s best response (BR) mapping is known to the leader. However, this is not always true in practice. In those cases the leader needs to simultaneously infer this BR function while fulfilling an internal objective. We study a setting in which the leader selects a control strategy that optimizes an objective given an initial belief about the follower’s best response. This belief is updated during the finite decision horizon, prompting the leader to reoptimize its control. We characterize the optimality guarantees of the SE solutions under this belief update for both open loop (OL) and feedback (FB) information structures. In particular, we show that it is possible that assuming an incorrect follower BR map obtains a lower cost over the game horizon than knowing the true BR. We support these claims with numerical examples in a linear quadratic (LQ) Stackelberg game.
@@ I Introduction
If the autonomous systems do not interact, i.e., if at most one autonomous system bases its strategy on the decisions of others, then that system can effectively treat the others as dynamic obstacles and perform obstacle avoidance by estimating their dynamics online. In this case, our research question becomes trivial: knowing the true model of the interaction leads better interaction outcomes [4]. However, autonomous systems inherently interact if they observe and avoid each other. Not modeling this interaction leads to incorrect outcome predictions and unsafe outcomes.
To address the posed question under autonomous interactions, this work models the interaction as a two player non-cooperative Stackelberg game [5, 6] where the leader estimates the follower’s BR and optimizes its own Stackelberg strategy under the current best BR guess simultaneously. We consider both OL and FB information structure with linear time invariant (LTI) dynamics and a finite and discrete time horizon. We focus on the problem from a strategy optimization perspective, and assume that the estimated follower’s BR comes from an oracle that may update the BR estimate during the control horizon.
Contributions. For OL Stackelberg equilibria, we prove a sufficient condition for when optimizing with respect to a correct BR leads to the optimal leader strategy (which implies that correct BR does not always lead to optimal leader strategy), and demonstrate its connection to the time inconsistency property of Stackelberg equilibria. For FB Stackelberg equilibria, we show that not all solutions imply that a correct BR leads to the optimal leader strategy despite its time consistency. We also prove that updating the leader’s strategies during the control horizon does not necessarily lead to better outcomes, even if we update with respect to a more accurate BR estimation. Finally, we demonstrate via numerical simulations of an LQ Stackelberg game with Bayesian updates of BR estimation that these theoretical results in fact frequently occur and lead to sub-optimal behavior when not accounted for.
@@ II Related Research
First introduced in [5] to model oligopolistic markets, the Stackelberg game and its dynamic extension [7, 8, 9] are used extensively in control theory to model lane-merging scenarios [10], pollution control [11], and power-grid optimization [12]. An important consideration in Stackelberg games is the information structure; common structures include open loop [13], feedback [14], and closed loop [15]. Both open loop and feedback Stackelberg equilibria have been analytically solved for the discrete-time linear quadratic (LQ) game [14], while [16] provides general optimality conditions for joint strategies. The closed loop Stackelberg equilibrium is more difficult to analytically derive in general. For LQ games, the two-step memory closed loop solutions are derived in [17, 15] based on a team-optimal strategy.
Although Stackelberg games assume that the leader knows the follower’s best response, this is not always true in practice. Estimating the opponents’ intentions is the focus of inverse learning [2] in dynamic game settings: [18] employs a Bayesian perspective to estimate opponent intentions, [19] uses inverse reinforcement learning [1] to model the interaction between autonomous cars and humans, and [20] provides a method to infer opponent objectives and compute a Nash equilibrium under noisy and partial observations. In particular, [18] derives an adaptive solution for finding Nash equilibria under partial observations while this paper explicitly addresses the optimality conditions of known Stackelberg equilibrium under such adaptive solutions. Finally, [21] studies a similar problem from an online learning perspective.
@@ III The Stackelberg game under unknown follower intentions
Notation We use ${\mathbb{R}}^{n}$ to denote a real-valued vector space of dimension $n$, ${\mathbb{R}}^{n\times m}$ to denote a real-valued matrix with dimensions $n\times m$, and $\mathbb{N}$ to denote the set of all natural numbers. The discrete number set $\{0,\ldots,T\}$ is denoted as $[T]$ and the discrete number set $\{t_{0},\ldots,t_{f}\}$ is denoted as $t_{0},t_{f}]$. Additionally, we denote the state trajectory as $\mathbf{x}=(x_{0},\ldots,x_{T})\in{\mathbb{R}}^{n(T+1)}$, the control trajectory as $\mathbf{u}=(u_{0},\ldots,u_{T-1})\in{\mathbb{R}}^{mT}$, and the strategy trajectory as $\bm{\pi}=(\pi_{0},\ldots,\pi_{T-1})\in{\mathbb{R}}^{(m\times n)T}$.
@@ III-A Stackelberg dynamic game
Under FB information structure, the control chosen is $u^{i}_{t}~=~-K^{i}_{t}x_{t}$ for $i\in\{L,F\}$. We use $\Pi^{i}_{0:T-1}$ to denote player $i$’s strategy domain over the finite horizon [T1][T-1], and $\bm{\pi}^{i}\in\Pi^{i}_{0:T-1}$ to denote a specific strategy for player $i$. When the information structure context $\mathcal{I}^{i}$ is clear, we use $\bm{\pi}^{i}$ to denote $\bm{\pi}^{i}(\mathcal{I}^{i})$.
@@
The follower’s objective $J^{F}_{0:T}$ is continuously differentiable and strictly convex in $\big(\mathbf{x},\mathbf{u}^{F}\big)$ for all $\mathbf{u}^{L}\in{\mathbb{R}}^{m_{L}T}$. The leader’s objective $J^{L}_{0:T}$ is continuously differentiable and strictly convex in $\big(\mathbf{x},\mathbf{u}^{L},\mathbf{u}^{F}\big)$.
Best response (BR). Under a leader’s strategy $\bm{\pi}^{L}$, the follower seeks to minimize its objective $J^{F}_{0:T}$ (5) using its strategy $\bm{\pi}^{F}$. This problem is formulated as
$\min_{\bm{\pi}^{F}\in\Pi^{F}_{0:T-1}}$ $\quad J^{F}_{0:T}\big(\mathbf{x},\bm{\pi}^{L}(\mathcal{I}^{L}),\bm{\pi}^{F}(\mathcal{I}^{F})\big)$ (6)
s.t. $\mathbf{x}=Hx_{0}+G^{L}\bm{\pi}^{L}+G^{F}\bm{\pi}^{F}.$
Under Assumption 1, the follower has a unique best response to any leader strategy $\bm{\pi}^{L}$ [6, Sec. 7.2] over the set of injective mappings from the leader’s strategy space to the follower’s, which we denote as
$\mathcal{F}=\{b\ |\ b:\Pi_{0:T-1}^{L}\mapsto\Pi_{0:T-1}^{F}\},$ (7)
and the follower’s BR is a mapping $b^{\star}\in\mathcal{F}$ such that $b^{\star}(\bm{\pi}^{L})$ is the solution to (6) for any $\bm{\pi}^{L}\in\Pi_{0:T-1}^{L}$. Next, we define the Stackelberg dynamic game.
@@
In a two player Stackelberg dynamic game with leader $L$ and follower $F$, the leader seeks a strategy $\bm{\pi}^{L}$ that solves
$\min_{\bm{\pi}^{L}\in\Pi^{L}_{0:T-1}}$ $J^{L}_{0:T}\big(\mathbf{x},\bm{\pi}^{L}(\mathcal{I}^{L}),b^{\star}(\bm{\pi}^{L})\big)$ (8)
s.t. $\mathbf{x}=Hx_{0}+G^{L}\bm{\pi}^{L}+G^{F}b^{\star}(\bm{\pi}^{L}),$
while the follower seeks a strategy $\bm{\pi}^{F}$ that solves (6). In particular, the leader always announces $\bm{\pi}^{L}$ first and the follower responds with $\bm{\pi}^{F}$.
Joint strategies that solve (8) under different information structure leads to different mathematical guarantees. We discuss each specific guarantee in later Sections IV-A and IV-B, respectively.
III-B Stackelberg game under follower BR updates
Under a known follower BR (6), we can find the SE for OL and FB information structures with guarantees [6]. To understand how simultaneous follower BR updates can affect these optimality concepts, we define a Stackelberg game with BR update in this section.
@@ III-B Stackelberg game under follower BR updates
Stackelberg game with BR update. Under Assumption 2, consider a Stackelberg game in which the leader does not know the follower’s true belief $b^{\star}$, but the leader first solves the $[T]$-horizon Stackelberg game (8) under belief $b^{1}$ to execute strategy $\bm{\pi}^{L1}\in\Pi^{L}_{0:T-1}$, the follower solves its $[T]$-horizon problem (6) to execute strategy $b^{\star}(\bm{\pi}^{L1})$, and the system dynamics evolve under (3) for horizon $0,\tau]$ as
@@
At the update time $\tau$, the leader solves an new optimal control problem with updated BR belief $b^{2}\in\mathcal{F}$ and shorter time horizon $\tau,T]$, given as
@@
$\text{s.t. }\quad x_{\tau:T}=H_{\tau:T}x_{\tau}+G^{L}_{\tau:T}\pi^{L}_{\tau:T-1}+G^{F}_{\tau:T}b^{2}(\pi^{L}_{\tau:T-1}).$
The leader computes and executes a new strategy under the new belief $b^{2}$, which we denote as $\hat{\pi}^{L2}_{\tau:T-1}$. Thus, the leader’s entire strategy over the horizon $[T]$ under BR update is
$\hat{\bm{\pi}}^{L}=\left[\pi^{L1}_{0:\tau-1},\hat{\pi}^{L2}_{\tau:T-1}\right].$ (12)
At $t=0$, the follower solves its $[T]$-horizon problem (6). When the leader updates its strategy to $\hat{\pi}^{L2}_{\tau:T-1}$ at $t=\tau$, the follower also update its best response to $b^{\star}(\hat{\pi}^{L2}_{\tau:T-1})$. Thus, the follower’s entire strategy over horizon $[T]$ is
$\hat{\bm{\pi}}^{F}=[\pi^{F1}_{0:\tau-1},b^{\star}(\hat{\pi}^{L2}_{\tau:T-1})],$ (13)
where $\pi^{F1}_{0:\tau-1}$ consists of the first $\tau$ elements of $b^{\star}(\bm{\pi}^{L1})$. Given the realized strategies $(\hat{\bm{\pi}}^{L},\hat{\bm{\pi}}^{F})$, the state trajectory in the Stackelberg game with BR update is given by
$\hat{\mathbf{x}}=[x^{1}_{0:\tau-1},x^{2}_{\tau:T}]\in{\mathbb{R}}^{n(T+1)},$ (14)
where $x^{1}_{0:\tau}$ is defined in (10) and $x^{2}_{\tau:T}$ is given by $x^{2}_{\tau:T}=H_{\tau:T}x^{1}_{\tau}+G^{L}_{\tau:T}\hat{\pi}^{L2}_{\tau:T-1}+G^{F}_{\tau:T}b^{\star}(\bm{\pi}^{L2}_{\tau:T-1})$ such that $x^{1}_{\tau}=x^{2}_{\tau}$.
To evaluate the optimality guarantees of the Stackelberg game with BR update, we compare the leader’s cost under the true BR $b^{\star}$ to the alternative beliefs $(b^{1},b^{2})\in\mathcal{F}$. To simplify the discussion, we break down the leader’s cost (5) into the components $J^{L}_{0:\tau-1}(b^{1})$ and $J^{L}_{{\tau:T}}(b^{2})$, defined as
$J^{L}_{0:\tau-1}(b^{1})$ $=\sum_{t=0}^{\tau-1}J^{i}_{t}\big(x^{1}_{t},\pi^{L1}_{t},\pi^{F1}_{t}\big),$ (15)
$J^{L}_{{\tau:T}}(b^{2})$ $=J_{T}^{i}(x^{2}_{T})+\sum_{t=\tau}^{T-1}J^{L}_{t}\big({x}^{2}_{t},\hat{{\pi}}_{t}^{L2},b^{\star}(\hat{{\pi}}_{t}^{L2})\big),$ (16)
such that $J^{L}_{0:T}(\hat{\mathbf{x}},\hat{\bm{\pi}}^{L},\hat{\bm{\pi}}^{F})=J^{L}_{0:\tau-1}(b^{1})+J^{L}_{{\tau:T}}(b^{2})$.
Having formulated the problem setting, we now state the central question addressed in this work.
@@ IV SE under BR Updates
This section addresses Problem 1 under OL and FB information structures (4). We first define the optimality conditions of both SE solutions, followed by a discussion on whether the SE solutions with BR updates and under the true BR belief satisfy (17).

### v2

@@ Abstract
We study a two-player dynamic Stackelberg game between a leader and a follower whose intention is unknown to the leader. Classical formulations of the Stackelberg equilibrium (SE) assume that the follower’s best response (BR) function is known to the leader. However, this is not always true in practice. We study a setting in which the leader receives updated beliefs about the follower BR before the end of the game, such that the update prompts the leader and subsequently the follower to re-optimize their strategies. We characterize the optimality guarantees of the SE solutions under this belief update for both open loop and feedback information structures. Interestingly, we prove that in general, assuming an incorrect follower’s BR can lead to more optimal leader costs over the entire game than knowing the true follower’s BR. We support these results with numerical examples in a linear quadratic (LQ) Stackelberg game, and use Monte Carlo simulations to show that the instances of incorrect BR achieving lower leader costs are non-trivial in collision avoidance LQ Stackelberg games.
@@ I Introduction
If the autonomous systems do not interact, i.e., if at most one autonomous system bases its strategy on the decisions of others, then that system can effectively treat the others as dynamic obstacles and perform obstacle avoidance by estimating their dynamics online. In this case, our research question becomes trivial: knowing the true model of the interaction leads to better interaction outcomes [4]. However, autonomous systems inherently interact if they observe and avoid each other. Not modeling this interaction leads to incorrect outcome predictions and unsafe outcomes.
To address the posed question under autonomous interactions, this work models the interaction as a two player non-cooperative Stackelberg game [5, 6] where the leader estimates the follower’s BR and optimizes its own Stackelberg strategy under the current best BR guess simultaneously. We consider both open loop (OL) and feedback (FB) information structure with linear time invariant (LTI) dynamics and a finite and discrete time horizon. We focus on the problem from a strategy optimization perspective, and assume that the estimated follower’s BR comes from an oracle that may update the BR estimate during the control horizon.
Contributions. For OL Stackelberg equilibria, we prove a sufficient condition for when optimizing with respect to a correct BR belief leads to the optimal leader strategy (which implies that the correct BR belief does not always lead to an optimal leader strategy), and demonstrate its connection to the time inconsistency property of Stackelberg equilibria. For FB Stackelberg equilibria, we show that not all solutions imply that a correct BR leads to the optimal leader strategy despite its time-consistency. Finally, we demonstrate via numerical simulations of an LQ Stackelberg game with Bayesian updates of BR estimation that these theoretical results in fact frequently occur and lead to sub-optimal behavior when not accounted for.
@@ II Related Research
First introduced in [5] to model oligopolistic markets, the Stackelberg game and its dynamic extension [7, 8, 9] are used extensively in control theory to model lane-merging scenarios [10], pollution control [11], and power-grid optimization [12]. An important consideration in Stackelberg games is the information structure; common structures include open loop [13], feedback [14], and closed loop [15]. Both open loop and feedback Stackelberg equilibria have been analytically solved for the discrete-time LQ game [14], while [16] provides general optimality conditions for joint strategies. The closed loop Stackelberg equilibrium is more difficult to analytically derive in general. For LQ games, the two-step memory closed loop solutions are derived in [17, 15] based on a team-optimal strategy.
Although Stackelberg games assume that the leader knows the follower’s BR, this is not always true in practice. Estimating the opponents’ intentions is the focus of inverse learning [2] in dynamic game settings: [18] employs a Bayesian perspective to estimate opponent intentions, [19] uses inverse reinforcement learning [1] to model the interaction between autonomous cars and humans, and [20] provides a method to infer opponent objectives and compute a Nash equilibrium under noisy and partial observations. In particular, [18] derives an adaptive solution for finding Nash equilibria under partial observations while this paper explicitly addresses the optimality conditions of known Stackelberg equilibrium under such adaptive solutions. Finally, [21] studies a similar problem from an online learning perspective.
@@ III The Stackelberg game under unknown follower intentions
Notation We use ${\mathbb{R}}^{n}$ to denote a real-valued vector space of dimension $n$, ${\mathbb{R}}^{n\times m}$ to denote a real-valued matrix with dimensions $n\times m$, and $\mathbb{N}$ to denote the set of all natural numbers. The discrete number set $\{0,\ldots,T\}$ is denoted as $[T]$ and the discrete number set $\{t_{0},\ldots,t_{f}\}$ is denoted as $t_{0},t_{f}]$. Additionally, we denote the state trajectory as $\mathbf{x}=(x_{0},\ldots,x_{T})\in{\mathbb{R}}^{n(T+1)}$, the control trajectory as $\mathbf{u}=(u_{0},\ldots,u_{T-1})\in{\mathbb{R}}^{mT}$, and the strategy trajectory as $\boldsymbol{\pi}=(\pi_{0},\ldots,\pi_{T-1})\in{\mathbb{R}}^{(m\times n)T}$.
@@ III-A Stackelberg dynamic game
Under FB information structure, the control chosen is $u^{i}_{t}~=~-K^{i}_{t}x_{t}$ for $i\in\{L,F\}$. We use $\Pi^{i}_{0:T-1}$ to denote player $i$’s strategy domain over the finite horizon [T1][T-1], and $\boldsymbol{\pi}^{i}\in\Pi^{i}_{0:T-1}$ to denote a specific strategy for player $i$. When the information structure context $\mathcal{I}^{i}$ is clear, we use $\boldsymbol{\pi}^{i}$ to denote $\boldsymbol{\pi}^{i}(\mathcal{I}^{i})$.
@@
The leader and follower objectives $J^{L}_{0:T}$ and $J^{F}_{0:T}$ are continuously differentiable and strictly convex in $\big(\mathbf{u}^{L},\mathbf{u}^{F}\big)$, and continuously differentiable and convex in $\mathbf{x}$.
Best response (BR). Under a leader’s strategy $\boldsymbol{\pi}^{L}$, the follower seeks to minimize its objective $J^{F}_{0:T}$ (5) using its strategy $\boldsymbol{\pi}^{F}$. This problem is formulated as
$\min_{\boldsymbol{\pi}^{F}\in\Pi^{F}_{0:T-1}}$ $\quad J^{F}_{0:T}\big(\mathbf{x},\boldsymbol{\pi}^{L}(\mathcal{I}^{L}),\boldsymbol{\pi}^{F}(\mathcal{I}^{F})\big)$ (6)
s.t. $\mathbf{x}=Hx_{0}+G^{L}\boldsymbol{\pi}^{L}+G^{F}\boldsymbol{\pi}^{F}.$
The follower’s BR is the $\mathop{\rm argmin}$ set of (6). Each leader strategy maps to a set of optimal follower responses. Under Assumption 1, this mapping is unique and continuous in the leader strategy $\boldsymbol{\pi}^{L}$ [6, Sec. 7.2]. We denote this continuous function as $b^{\star}:\Pi_{0:T-1}^{L}\mapsto\Pi_{0:T-1}^{F}$, so that $b^{\star}(\boldsymbol{\pi}^{L})$ is the optimal follower strategy under (6) for all $\boldsymbol{\pi}^{L}\in\Pi_{0:T-1}^{L}$. Additionally, we use $\mathcal{F}$ to denote the set of all continuous follower response functions, given by
$\mathcal{F}=\{b\ |\ b:\Pi_{0:T-1}^{L}\mapsto\Pi_{0:T-1}^{F}\}.$ (7)
Next, we define the Stackelberg dynamic game.
@@
In a two player Stackelberg dynamic game with leader $L$ and follower $F$, the leader seeks a strategy $\boldsymbol{\pi}^{L}$ that solves
$\min_{\boldsymbol{\pi}^{L}\in\Pi^{L}_{0:T-1}}$ $J^{L}_{0:T}\big(\mathbf{x},\boldsymbol{\pi}^{L}(\mathcal{I}^{L}),b^{\star}(\boldsymbol{\pi}^{L})\big)$ (8)
s.t. $\mathbf{x}=Hx_{0}+G^{L}\boldsymbol{\pi}^{L}+G^{F}b^{\star}(\boldsymbol{\pi}^{L}),$
while the follower seeks a strategy $\boldsymbol{\pi}^{F}$ that solves (6). In particular, the leader always announces $\boldsymbol{\pi}^{L}$ first and the follower responds with $\boldsymbol{\pi}^{F}$.
Under OL and FB information structures and with known follower’s BR, joint strategies that solve (8) exist and lead to different mathematical guarantees [6]. We discuss each specific guarantee in later Sections IV-A and IV-B, respectively.
III-B Stackelberg game under follower BR belief updates
In this subsection, we formulate a Stackelberg dynamic game in which the leader does not know the follower’s BR $b^{\star}$ but has time-varying beliefs about it.
@@ III-B Stackelberg game under follower BR belief updates
Stackelberg game with BR belief update. Under Assumption 2, consider a Stackelberg game in which the leader does not know the follower’s true BR $b^{\star}\in\mathcal{F}$. In this setting, the leader first solves the $[T]$-horizon Stackelberg game (8) under belief $b^{1}\in\mathcal{F}$ to compute a strategy $\boldsymbol{\pi}^{L1}\in\Pi^{L}_{0:T-1}$ and the follower solves its $[T]$-horizon problem (6) under its true BR $b^{\star}$ to compute a strategy $b^{\star}(\boldsymbol{\pi}^{L1})$. For $\tau\in[T]$ time steps, the system dynamics (3) evolve under these strategies as
@@
At the update time $\tau$, the system state arrives at $x_{\tau}$. The leader receives an updated BR belief $b^{2}\in\mathcal{F}$, re-solves the Stackelberg game under $b^{2}$ for the remaining horizon $\tau,T]$, and executes a new optimal strategy $\hat{\pi}^{L2}_{\tau:T-1}$ that minimizes
@@
$\text{s.t. }\quad x_{\tau:T}=H_{\tau:T}x^{1}_{\tau}+G^{L}_{\tau:T}\pi^{L}_{\tau:T-1}+G^{F}_{\tau:T}b^{2}(\pi^{L}_{\tau:T-1}).$
Once the leader announces new strategy $\hat{\pi}^{L2}_{\tau:T-1}$, the follower also updates its best response to $b^{\star}(\hat{\pi}^{L2}_{\tau:T-1})$, and the system dynamics evolve from $x_{\tau}$ as
$x^{2}_{\tau:T}=H_{\tau:T}x^{1}_{\tau}+G^{L}_{\tau:T}\hat{\pi}^{L2}_{\tau:T-1}+G^{F}_{\tau:T}b^{\star}(\hat{\pi}^{L2}_{\tau:T-1}).$ (12)
Over the time horizon $[T]$ and under the BR belief update, the leader’s composite strategy is given by
$\hat{\boldsymbol{\pi}}^{L}=\left[\pi^{L1}_{0:\tau-1},\hat{\pi}^{L2}_{\tau:T-1}\right],$ (13)
the follower’s composite strategy given by
$\hat{\boldsymbol{\pi}}^{F}=[\pi^{F1}_{0:\tau-1},b^{\star}(\hat{\pi}^{L2}_{\tau:T-1})],$ (14)
and the composite state trajectory is given by
$\hat{\mathbf{x}}=[x^{1}_{0:\tau-1},x^{2}_{\tau:T}]\in{\mathbb{R}}^{n(T+1)}.$ (15)
To highlight how the leader’s total cost depends on the change in its BR belief, we break down the leader’s cost (5) into the components $J^{L}_{0:\tau-1}(b^{1})$ and $J^{L}_{{\tau:T}}(b^{2})$, defined as
$J^{L}_{0:\tau-1}(b^{1})$ $=\sum_{t=0}^{\tau-1}J^{i}_{t}\big(x^{1}_{t},\pi^{L1}_{t},b^{\star}({{\pi}}_{t}^{L1})\big),$ (16)
$J^{L}_{{\tau:T}}(b^{2})$ $=J_{T}^{i}(x^{2}_{T})+\sum_{t=\tau}^{T-1}J^{L}_{t}\big({x}^{2}_{t},\hat{{\pi}}_{t}^{L2},b^{\star}(\hat{{\pi}}_{t}^{L2})\big).$ (17)
such that $J^{L}_{0:T}(\hat{\mathbf{x}},\hat{\boldsymbol{\pi}}^{L},\hat{\boldsymbol{\pi}}^{F})=J^{L}_{0:\tau-1}(b^{1})+J^{L}_{{\tau:T}}(b^{2})$.
This BR belief update setting captures applications of Stackelberg games in iterative or data-driven settings such as [23, 24]. However, it differs from inverse game-theoretic approaches in which the control is strictly selected to uncover follower BR [25, 26].
The key question we explore in this manuscript is whether the leader’s cost is the lowest when it knows the true BR $b^{\star}$ versus all possible BR beliefs $(b^{1},b^{2})\in\mathcal{F}$.
@@ IV SE under BR Updates
This section addresses Problem 1 under OL and FB information structures (4). We first define the optimality conditions of both SE solutions, followed by a discussion on whether the SE solutions with BR belief updates under the true BR belief satisfy (18).